# Berezinskii-Kosterlitz-Thouless transition in the XY model on the honeycomb lattice: A comprehensive Monte Carlo analysis

### v2

@@ IV Results
Figure 1 illustrates the honeycomb lattice and its transposition to the equivalent square lattice used in the simulations. This transposition significantly simplifies the implementation of the numerical algorithm in the Monte Carlo simulations without any loss of generality.
All simulations were conducted on lattices with sizes ranging from $L=6$ to $L=128$ with periodic boundary conditions. For the simulated annealing (SA) protocol, we set the initial temperature to $\beta_{0}=0.1$ and the final temperature to $\beta_{n}=3$, with $n=100$ and $m_{s}=10^{6}$. This involved a total of $10^{8}$ Monte Carlo (MC) sweeps for each temperature. Additionally, ten independent simulations were performed for each lattice size to compute statistical errors. For the largest lattice size, all spin configurations were saved at each temperature to analyze the structural order of the spins and calculate the vortex density.
@@
In the Wang-Landau (WL) procedure, we adopted a similar approach to that used in Ref. Xu and Ma (2007) for calculating the density of states in continuous models. Since the boundaries of the density of states spectrum exhibit large fluctuations, the flatness criterion requires a significant amount of computational time. To address this issue, we considered the entire energy range, which is $3N$ (where $N=L\times L$), and divided it into $N^{2}/4$ energy levels (bins). The standard flatness criterion is applied only to the middle spectrum range, specifically $[-N^{2}/4+N,N^{2}/4-N]$. For the boundary levels, we used an 80% flatness criterion. This approach significantly reduces computational time.
Figure 2 presents various thermodynamic quantities as a function of temperature for both the SA (indicated by dots) and WL (represented by solid lines) procedures. Given that the WL simulation is limited to smaller system sizes, we only present results for $L$ values up to $32$ in this case. As shown, the specific heat approaches $1/2$ as $T\to 0$, regardless of the lattice size, which aligns with the expectations of the equipartition theorem. Although both the specific heat and susceptibility exhibit a peak slightly above $T=0.6$, these peaks are not used to determine the BKT transition temperature due to their lack of accuracy. Additionally, the peak of the specific heat does not display a consistent behavior as the system size increases. However, a finite-size scaling (FSS) analysis of the critical temperature using this data yields values that are lower than the expected theoretical value of $1/\sqrt{2}$, as summarized in Table 1.
@@
Some authors use the Binder cumulant of the order parameter crossing technique together with another sophisticated numerical procedure to determine the critical temperatureSilva et al. (2022). Figure 2 also presents the behavior of this quantity as a function of temperature for several lattice sizes. As expected, the Binder cumulants approach a value of $1/3$ in the disordered (high-temperature) state. While for the ordered (low temperature) state they approach the value of $2/3$.
@@
For a more precise determination of the critical temperature, we use the helicity modulus obtained from the simulated annealing (SA) procedure as a function of the inverse temperature for various lattice sizes, as illustrated in Fig. 3. We are using inverse temperature here because the SA results depend on $\beta$, and this choice helps to avoid rounding errors during the finite-size scaling analysis. The theoretical curve, represented by $2/\pi\beta$, is also shown. The discontinuity of the helicity modulus near the transition is not very pronounced. As a discontinuity occurs in the thermodynamic limit, we can use the intersection of the theoretical curve $2/\pi\beta$ with the simulated values to estimate the critical temperature. As the lattice size increases, this intersection point shifts to the right, approaching the critical temperature value in the thermodynamic limit. One challenge we face is accurately determining the intersection point, since the data obtained from SA are not continuous. Additionally, achieving a larger sample size with SA would entail significant computational costs. This issue does not arise with WL simulations, but they are limited as system sizes increase.
@@
Based on this, we individually calculated the values of $\beta_{BKT}$ for each lattice and performed a finite-size scaling analysis, as shown in Fig. 5, for both $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$. Using the helicity modulus, we obtained $\beta_{BKT}=1.685\pm 0.003$. Note that this value is much higher than the theoretical $\sqrt{2}$. Using the WL technique, we obtained $\beta_{BKT}=1.627\pm 0.004$. The difference is due to the maximum limited size of $L=32$ used in the latter. But this value is still higher than the theoretical one.
Obtaining the critical temperature through $\langle\Upsilon_{4}\rangle$ is straightforward. Thus, a value of $\beta_{c}$ was obtained for each lattice size and, again, we performed the FSS to calculate $\beta_{BKT}$, as shown in Fig. 5. In this case, we obtained $\beta_{BKT}=1.627\pm 0.012$ from SA and $\beta_{BKT}=1.648\pm 0.007$ from WL. Here one can see a little discrepancy between the outcomes from the two technique and this is because the WL allows for a high resolution search for the critical temperature corresponding to the $\langle\Upsilon_{4}\rangle$. But still, these results confirm that the theoretical prediction can not be supported by neither numerical analysis.
@@
Another powerful method for analyzing critical properties and determining the critical temperature is the crossing of the Binder cumulant. Figure 6 illustrates the cumulant crossings for different system sizes as a function of the inverse temperature obtained through SA. Despite the apparent randomness of the crossings, it is evident that the cumulant exhibits scaling behavior. Albuquerque et al. employed a sophisticated extrapolation scheme of the crossings to obtain the critical temperature of a modified XY model on the square latticeSilva et al. (2022). However, since we are primarily interested in the general aspects of the cumulant crossings to support our findings, we will use a method similar to that described in Ref. Liu et al. (2018). Figure 6 (bottom) also shows how the inverse temperature at which the crossings occur varies with lattice sizes. A rough estimate of the inverse temperature in the thermodynamic limit gives $\beta_{cross}=1.724\pm 0.002$. Notably, this value is even higher than those obtained from the helicity modulus compared to the theoretical prediction. Very recently, F. Jiang reported neural network calculations of the same problemJiang (2024) and found a value of $\approx 1.75$ for $\beta_{BKT}$.
It is helpful to illustrate the topology of spin configurations at different temperatures as the BKT transition passes through. Figure 7 displays six configurations for the lattice size of $L=128$. The color pattern indicates the individual orientation of each spin and follows the following scale (in radians): dark red – $+\pi/2$ rad; dark blue – $-\pi/2$ rad; white – 00 rad; black – ±$\pm 2\pi$ rad. For the temperature $T=0.25$, below the critical temperature, we can see the formation of a pattern in the orientation of the spins (predominance of spins oriented “vertically”) but which is not uniform throughout the lattice; this occurs due to the presence of vortex-antivortex pairs. For $T=0.57$, in the region where the transition occurs, it is easy to see how the system becomes quite heterogeneous, possibly due to the dissociation of the vortex-antivortex pairs. As the temperature rises above $T_{BKT}$, the spins become misaligned, leading to an increase in disorder until complete dissociation occurs. At this point, the system reaches a saturation point in the number of vortices.
@@
To perform a quantitative analysis of the dissociation of vortex-antivortex pairs, we can calculate the vortex density. A vortex is identified when the sum of $\theta_{i}-\theta_{j}$ along a closed path at each lattice site is a multiple of ±$\pm 2\pi$. Using this criterion, we determine the vortex density for various temperatures within our simulated range on a lattice of size $L=128$, where the vortex density is defined as the number of vortices divided by $L^{2}$. Figure 8 illustrates how the vortex density varies as a function of temperature. As anticipated, the density increases with higher temperaturesSun et al. (2022). It is also noteworthy that the rate of increase in density is more pronounced near the BKT transition temperature, which occurs at a value significantly lower than the theoretical prediction. This observation supports our finite-size scaling (FSS) analysis.
@@
where $2\mu$ is the angular coefficient of this function. Using this relationship, we adjusted the vortex density values close to the critical temperature and obtained $2\mu=4.0\pm 0.5$ in units of $J$. This value is less than half of what was obtained for the square latticeGupta and Baillie (1992). This result reinforces the idea of the instability of the honeycomb lattice.
@@ V Conclusions
We carefully analyzed the XY model on the honeycomb lattice using the numerical Monte Carlo method, applying two techniques: the Simulated Annealing (SA) protocol and entropic simulations based on the Wang-Landau (WL) algorithm. Our primary objective was to calculate the critical transition temperature and to compare our findings with theoretical predictions. To achieve this, we used the second- and fourth-order helicity modulus as order parameters, as these properties exhibit a discontinuity during the phase transition. Our best estimates were obtained using the SA protocol, which provided values for the inverse transition temperature of $\beta_{BKT}=1.627(12)$ (based on fourth-order helicity) and $T_{BKT}=1.685(3)$ (based on second-order helicity). Notably, these values are significantly higher than the analytical prediction of $\sqrt{2}$. By analyzing the crossing of the Binder cumulant, we obtained an even higher value of $\beta_{cross}=1.724\pm 0.002$. This value is closer to the results obtained through neural network calculations. Additionally, we calculated the vortex density, which allowed us to estimate the formation energy of a pair of vortices. The obtained value was $2\mu=4.0\pm 0.5$, in units of $J$. This is more than 50% smaller than that of the square lattice.

### v3

@@ IV Results
Figure 1 illustrates the honeycomb lattice and its transposition to the equivalent square lattice used in the simulations. This transposition significantly simplifies the implementation of the numerical algorithm in the MC simulations without any loss of generality.
All simulations were conducted on lattices with sizes ranging from $L=6$ to $L=128$ with periodic boundary conditions. For the SA protocol, we set the initial temperature to $\beta_{0}=0.1$ and the final temperature to $\beta_{n}=3$, with $n=100$ and $m_{s}=10^{6}$. This involved a total of $10^{8}$ MC sweeps for each temperature. Additionally, ten independent simulations were performed for each lattice size to compute statistical errors. For the largest lattice size, all spin configurations were saved at each temperature to analyze the structural order of the spins and calculate the vortex density.
@@
In the WL procedure, we adopted a similar approach to that used in Ref. Xu and Ma (2007) for calculating the density of states in continuous models. Since the boundaries of the density of states spectrum exhibit large fluctuations, the flatness criterion requires a significant amount of computational time. To manage this limitation, we considered the entire energy range, which is $3N$ (where $N=L\times L$), and divided it into $N^{2}/4$ energy levels (bins). The standard flatness criterion is applied only to the middle spectrum range, specifically $[-N^{2}/4+N,N^{2}/4-N]$. For the boundary levels, we used an 80% flatness criterion. This approach significantly reduces computational time.
Figure 2 presents various thermodynamic quantities as a function of temperature for both SA (indicated by dots) and WL (represented by solid lines) procedures. Since the WL simulation is constrained to smaller system sizes, we present results only for $L$ values up to 32. As shown, the specific heat approaches $1/2$ as $T\to 0$, regardless the lattice size, which aligns with the expectations of the equipartition theorem. Although both the specific heat and susceptibility exhibit a peak slightly above $T=0.6$, these peaks are not used to determine the BKT transition temperature due to their lack of accuracy. Additionally, the peak of the specific heat does not display a consistent behavior as the system size increases. However, a FSS analysis of the critical temperature using this data yields values that are lower than the expected theoretical value of $1/\sqrt{2}$, as summarized in Table 1.
@@
Some authors use the Binder cumulant of the order parameter crossing technique together with another sophisticated numerical procedure to determine the critical temperatureSilva et al. (2022). Figure 2 also presents the behavior of this quantity as a function of temperature for several lattice sizes. As expected, the Binder cumulants approach a value of $1/3$ in the disordered (high-temperature) state. Although for the ordered (low temperature) state they approach the value of $2/3$.
@@
For a more precise determination of the critical temperature, we use the helicity modulus obtained from the SA procedure as a function of the inverse temperature for various lattice sizes, as illustrated in Fig. 3. We use inverse temperature in this context because the SA results depend on $\beta$. This choice also minimizes rounding errors during the finite-size scaling analysis. The theoretical curve, represented by $2/\pi\beta$, is also shown. The discontinuity of the helicity modulus near the transition is not very pronounced. As a discontinuity occurs in the thermodynamic limit, we can use the intersection of the theoretical curve $2/\pi\beta$ with the simulated values to estimate the critical temperature. As the lattice size increases, the intersection point shifts to the right, converging toward the critical temperature in the thermodynamic limit. A key challenge lies in accurately identifying this intersection, as the SA data are discrete. Moreover, obtaining a larger sample size with SA would incur significant computational costs. This issue does not arise with WL simulations, but they are limited as system size increases.
@@
Based on this, we individually calculated the values of $\beta_{BKT}$ for each lattice and performed a finite-size scaling analysis, as shown in Fig. 5, for both $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$. Using the helicity modulus, we obtained $\beta_{BKT}=1.687\pm 0.003$. It is important to note that this value is significantly higher than the theoretical prediction of $\sqrt{2}$. Although the WL technique yields $\beta_{BKT}=1.627\pm 0.004$. The difference is due to the maximum limited size of $L=32$ used in the latter. But this value is still higher than the theoretical one.
Obtaining the critical temperature through $\langle\Upsilon_{4}\rangle$ is relatively straightforward. We calculated $\beta_{c}$ for each lattice size and performed the FSS to determine $\beta_{BKT}$, as shown in Fig. 5. We found $\beta_{BKT}=1.635\pm 0.011$ using SA and $\beta_{BKT}=1.648\pm 0.007$ using the WL method. There is a slight discrepancy between the results from the two approaches because the WL method allows a higher-resolution search for the critical temperature corresponding to $\langle\Upsilon_{4}\rangle$. Additionally, it is important to note that the values obtained from $\langle\Upsilon\rangle$ and $\langle\Upsilon_{4}\rangle$ do not agree even within their error margins. This discrepancy may be due to the gradual approach of the minimum of $\langle\Upsilon_{4}\rangle$ towards the critical temperature in the thermodynamic limit, suggesting that the scaling relation given by Eq. 22 may not be appropriate in this context. Nevertheless, these results indicate that the theoretical predictions are not supported by numerical analysis.
@@
An alternative yet powerful approach for identifying critical properties and determining the transition temperature is the analysis of the crossing points of the Binder cumulant, a technique widely used in FSS studiesBinder (1981). Figure 6 illustrates the cumulant crossings for different system sizes as a function of the inverse temperature obtained through SA. Despite the apparent randomness of the crossings, it is evident that the cumulant exhibits scaling behavior. Albuquerque et al. employed a sophisticated extrapolation scheme of the crossings to obtain the critical temperature of a modified XY model on the square latticeSilva et al. (2022). However, since we are primarily interested in the general aspects of the cumulant crossings to support our findings, we will use a method similar to that described in Ref. Liu et al. (2018). Figure 6 (bottom) also shows how the inverse temperature at which the crossings occur varies with lattice sizes. A rough estimate of the inverse temperature in the thermodynamic limit gives $\beta_{cross}=1.724\pm 0.002$. Notably, this value is even higher than those obtained from the helicity modulus compared to the theoretical prediction. Very recently, F. Jiang reported neural network calculations of the same problemJiang (2024) and found a value of $\approx 1.75$ for $\beta_{BKT}$, which further supports our findings.
It is helpful to illustrate the topology of spin configurations at different temperatures as the BKT transition passes through. Figure 7 displays six configurations for the lattice size of $L=128$. The color pattern indicates the individual orientation of each spin and follows the following scale (in radians): dark red – $+\pi/2$ rad; dark blue – $-\pi/2$ rad; white – 00 rad; black – ±$\pm 2\pi$ rad. For the temperature $T=0.25$, below the critical temperature, we can see the formation of a pattern in the orientation of the spins (predominance of spins oriented “vertically”) but it is not uniform throughout the lattice; this occurs due to the presence of vortex-antivortex pairs. For $T=0.57$, in the region where the transition occurs, it is easy to see how the system becomes quite heterogeneous, possibly due to the dissociation of the vortex-antivortex pairs. As the temperature rises above $T_{BKT}$, the spins become misaligned, leading to an increase in disorder until complete dissociation occurs. At this point, the system reaches a saturation point in the number of vortices.
@@
To perform a quantitative analysis of the dissociation of vortex-antivortex pairs, we can calculate the vortex density. A vortex is identified when the sum of $\theta_{i}-\theta_{j}$ along a closed path at each lattice site is a multiple of ±$\pm 2\pi$. Using this criterion, we determine the vortex density for various temperatures within our simulated range on a lattice of size $L=128$, where the vortex density is defined as the number of vortices divided by $L^{2}$. Figure 8 illustrates how the vortex density varies as a function of temperature. As anticipated, the density increases with higher temperaturesSun et al. (2022). It is also noteworthy that the rate of increase in density is more pronounced near the BKT transition temperature, which occurs at a value significantly lower than the theoretical prediction. This observation supports our FSS analysis.
@@
where $2\mu$ is the angular coefficient of this function. Using this relationship, we adjusted the vortex density values close to the critical temperature and obtained $2\mu=4.0\pm 0.5$ in units of $J$. This value is less than half of what was obtained for the square latticeGupta and Baillie (1992). This result is justified by the idea of the instability of the honeycomb lattice to support the spin long-range order, as investigated by Wojtkiewicz et al.Wojtkiewicz et al. (2023).
@@ V Conclusions
We carefully analyzed the XY model on the honeycomb lattice using the numerical Monte Carlo method, applying two techniques: the Simulated Annealing protocol and entropic simulations based on the Wang-Landau algorithm. Our primary objective was to calculate the critical transition temperature and to compare our findings with theoretical predictions. To achieve this, we used the second- and fourth-order helicity modulus as order parameters, as these properties exhibit a discontinuity during the phase transition. Our best estimates were obtained using the SA protocol, which provided values for the inverse transition temperature of $\beta_{BKT}=1.635(11)$ (based on fourth-order helicity) and $T_{BKT}=1.687(3)$ (based on second-order helicity). Notably, these values are significantly higher than the analytical prediction of $\sqrt{2}$. By analyzing the crossing of the Binder cumulant, we obtained an even higher value of $\beta_{cross}=1.724\pm 0.002$. This value is closer to the results obtained through neural network calculations. Additionally, we calculated the vortex density, which allowed us to estimate the formation energy of a pair of vortices. The obtained value was $2\mu=4.0\pm 0.5$, in units of $J$. This is more than 50% smaller than that of the square lattice.